OpenStudy (anonymous):

Hello. I have an issue with factoring. The problem is 64x^3-27. I think this is unfactorable. Am I right or wrong?

OpenStudy (anonymous):

Try the difference of two cubes formula:\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (unklerhaukus):

\[64=4^3\]\[27=3^3\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Thank you both. One answer showed me where to look in my book and the other gave me the idea of what to do.